OpenStudy (anonymous):

a rectangular sheet of perimeter 27 cm and dimensions x cm by y cm is to be rolled into a cylinder. what values of x and y give the largest volume?

OpenStudy (anonymous):

me first buhahaha

OpenStudy (anonymous):

x = 2*pi*r r = x/2pi

OpenStudy (anonymous):

V = 1/(4pi) (27/2 * x^2 - x^3) dV/dx = 1/(4pi) (27/2 * 2x - 3x^2) dV/dx = 1/(4pi) (27x - 3x^2) now, take dV/dx = 0 1/(4pi) (27x - 3x^2) = 0 multiply by 4pi to both sides, it can be 27x - 3x^2 = 0 then divide by -3, giving us x^2 - 9x = 0 factor out! x(x-9) = 0 the zeroes will satisfies only x = 9